OpenStudy (anonymous):

For each formula, describe the four quantities. Then explain how you can find the fourth quantity if you know the values of the other three. A) an=a1 + (n-1) d B) sn= n/2 ( a1-an) C) an=a1r ^ n-1 D) Sn = a1 ( 1-r^n) / 1-r

OpenStudy (anonymous):

In an Arithmetic Sequence the difference between one term and the next is a constant. A1 is the first term, d is the common difference between each and every number after the first term and n is the number of terms in which one would plug in 4. a_n = a_1 + (n – 1)d This is the formula of the arithmetic sequence, in an arithmetic sequence it is going from one number by a common difference, when we sum up a number of terms, we first take the number of terms multiplied by the first term + the last term divided by 2. S_n=n(a_1+a_n )/2 In a geometric series, there is a series where each successive term is the previous term multiplied by a fixed value, common ratio (r). 〖a_1 r〗^(1-1),〖a_1 r〗^(2-1),〖a_1 r〗^(3-1),〖a_1 r〗^(4-1) a_n=a_1 r^(n-1) We can then find the sum of the geometric series using this formula. Which is done by distributing the term “a” to 1 minus the common ratio with a power containing the number of terms divided by 1 minus the common ratio. S_n=a(1-r^n )/((1-r))

OpenStudy (anonymous):

I'm just wondering if I described the four quantities right and if I'm supposed to solve for n only.

OpenStudy (anonymous):

@zepdrix , @mathstudent55 , @ganeshie8

ganeshie8 (ganeshie8):

Overall it looks good, I see one mistake : whats the difference between a `sequence` and a `series` ?

ganeshie8 (ganeshie8):

You're using both the terms as synonyms... but they mean different things... do you know hte difference between the two ? :)

OpenStudy (anonymous):

a series is a sum of terms and a sequence is a list of numbers with some form of pattern.

OpenStudy (anonymous):

But may I know where I made the small mistake?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@ganeshie8 , would you be able to show me where you had spotted the mistake?

ganeshie8 (ganeshie8):

exactly ! below is a sequence : \[\large 1,~2,~3,~4,~\cdots\] below is a series: \[\large 1+2+3+4+\cdots\]

ganeshie8 (ganeshie8):

A) an=a1 + (n-1) d `arith metic sequence formula` B) sn= n/2 ( a1-an) `arithmetic series formula` C) an=a1r ^ n-1 `geometric sequence formula` D) Sn = a1 ( 1-r^n) / 1-r `geometric series formula`

ganeshie8 (ganeshie8):

`Sn` is a series formula, `an` is a sequence formula... okay ?

OpenStudy (anonymous):

okay I will make sure to do that and it says "find the fourth quantity", but im not quite sure what its talking about.

ganeshie8 (ganeshie8):

look at the arithmetic sequence formula : \[\large a_n = a_1 + (n – 1)d \] how many variables(quantities) do you see ?

OpenStudy (anonymous):

4

ganeshie8 (ganeshie8):

and suppose if i tell you ANY 3 of those variables, will you be able to solve for the other variable ?

OpenStudy (anonymous):

an, a1, n, and d

ganeshie8 (ganeshie8):

yes!

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good :) tell me how you will solve the other variable ?

ganeshie8 (ganeshie8):

just describe, this is more of a writing problem...

OpenStudy (anonymous):

so if i were to solve for a1 I would first divide dn on both sides and add d on both sides, which would leave me with a1=d+ n(a-d)

ganeshie8 (ganeshie8):

kindof.. \[\large a_n = a_1 + (n – 1)d\] to solve \(a_1\), you would simply subtract \((n-1)d\) both sides : \[\large a_n - (n – 1)d = a_1 \]

OpenStudy (anonymous):

so i don't distribute then, i see.

ganeshie8 (ganeshie8):

yees we don't need to distribute cuz \(a_1\) was isolated already...

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

just add this step to the other 3 formulas as well, good luck !

OpenStudy (anonymous):

alright thanks for the help